OpenStudy (anonymous):

Let f(x)=-5x-4 and g(x)=6x-7. Find f(x)+g(x).

OpenStudy (anonymous):

f(x)+g(x)=-5x+4+6x-7 =x-3 do you understand??

OpenStudy (anonymous):

yea, you subtracted 6x and 5x then 4 from 7 right?

OpenStudy (anonymous):

yes i did.

OpenStudy (anonymous):

the answer is not =x-3

OpenStudy (anonymous):

=x-11 is true answer

OpenStudy (anonymous):

i had a mistake. so sorry. answer is x-11. thank you @baziham

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

nevermind i see it. thanks

OpenStudy (anonymous):

f(x)=-5x-4 g(x)=6x-7 ------------- 1x-11